if A= [ 1 -3 2 and B= [ 2 1 5
4 1 -1 -1 -2 -2
-3 2 5 ] 3 1 2 ]

then find (AB)^-1
1
Expert's answer
2020-12-20T18:54:18-0500

Given matrices are

"A = \begin{bmatrix}
 1 & -3 & 2 \\
 4 & 1 & -1 \\
 3 & 2 & 5 
\end{bmatrix}" and "B = \begin{bmatrix}
 2 & 1 & 5 \\
 -1 & -2 & -2 \\
 3 & 1 & 2
\end{bmatrix}"



Now, "AB = \begin{bmatrix}
 1 & -3 & 2 \\
 4 & 1 & -1 \\
 3 & 2 & 5
\end{bmatrix} \begin{bmatrix}
 2 & 1 & 5 \\
 -1 & -2 & -2 \\
 3 & 1 & 2
\end{bmatrix} = \begin{bmatrix}
 11 & 9 & 15 \\
 4 & 1 & 16 \\
 19 & 4 & 21
\end{bmatrix}"


Determinant of the matrix AB is 1471.

Co-factors are:

"a_{11} = (21-64) = -43"

"a_{12}=-(84-304)=220"

"a_{13}= (16-19) = -3"

"a_{21}=-(189-60)=-129"

"a_{22}=(231-285)=-54"

"a_{23}=-(44-171)=127"

"a_{31}=(144-15)=129"

"a_{32}=-(176-60)=-116"

"a_{33}=(11-36)=-25"



"adj (AB) = \begin{bmatrix}
 -43 & -129 & 129 \\
 220 & -54 & -116 \\
 -3 & 127 & -25
\end{bmatrix}"

"adj (AB) ="

Then inverse of the matrix AB is,

"(AB)^{-1} = \frac{1}{1471}\begin{bmatrix}
 -43 & -129 & 129 \\
 220 & -54 & -116 \\
 -3 & 127 & -25
\end{bmatrix}"